OpenStudy (anonymous):

How would you graph this? f(x)=3^(|x|+2)

OpenStudy (anonymous):

Like this

OpenStudy (anonymous):

Thanks for the graph, but how can you graph it yourself? using what calculations?

OpenStudy (anonymous):

Just make a T table. With x on one side of the T and y on the other side of the T. So if you pick x=0 then y=3^(0+2) or 9 Select another value for x, like 1. The absolute value of 1 is 1. So you now have y=3^(1+2) or 3^3=27 Select another value for x, like -1. The absolute value of -1 is 1. So you now have y=3^(1+2) or 3^3=27 Select enough points until you get the general shape.